If cos B = -8/17 in quadrant II, find sin 7B.
Answer by lwsshak3(11628)

If cos B = -8/17 in quadrant II, find sin 7B
**
arccosB(-8/17)≈118.07º
7B≈7*118.07=826.51º
sin7B=sin(826.51º)
reference angle≈826.51-720≈106.51 (quadrant II were sin>0)
sin 7B≈.96